A laser weld detachedretains emits light with awavelength 652nm.in pulses that are of 20m.duration . The average power during each pulse is 0.6w. The energy each pulse and single photon are

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
E(ev) = 1240/652 = 1.9 ev (in a single photon)
Energy in 20 ms = 20 * 0.6 * 10^-3 = 1.2 * 10^-2 joules
= 1.2 * 10^-2/ 1.6 * 10^-19
= 7.5 * 10^16 ev
 
 
 
Regards
Arun (askIITians forum expert)